OpenStudy (anonymous):

Can someone check my answers on any of these problems? Simplify the expression where possible. (-5x^2)^3 I think that the answer is 75x^2 Simplify the expression where possible. (3b^3)^4 I had 3b^12, but that was wrong. Simplify the expression where possible. (-4x^2)^2 I have -4x^4

OpenStudy (anonymous):

oh, ok

OpenStudy (anonymous):

Any advice for the others?

OpenStudy (anonymous):

first one should be \(125x^6\)

OpenStudy (anonymous):

\[(3b^3)^4=3^4x^{3\times 4} =81x^{12}\]

OpenStudy (anonymous):

oops i meant \(81b^{12}\)

OpenStudy (anonymous):

Okay,

OpenStudy (anonymous):

\[(-4x^2)^2 =(-4)^2\times (x^2)^2=16x^4\]

OpenStudy (anonymous):

Thanks so much!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

@satellite73 the first one was considered wrong... any other ideas?

OpenStudy (anonymous):

yes, because \(-5)^2=-5\times -5\times -5=-125\)

OpenStudy (anonymous):

try \(-125x^6\)

OpenStudy (anonymous):

okay...

OpenStudy (anonymous):

Awesome, it worked!